# What Matters in Hierarchical Search for Combinatorial Reasoning Problems?

### v1

@@ Abstract
Efficiently tackling combinatorial reasoning problems, particularly the notorious NP-hard tasks, remains a significant challenge for AI research. Recent efforts have sought to enhance planning by incorporating hierarchical high-level search strategies, known as subgoal methods. While promising, their performance against traditional low-level planners is inconsistent, raising questions about their application contexts. In this study, we conduct an in-depth exploration of subgoal-planning methods for combinatorial reasoning. We identify the attributes pivotal for leveraging the advantages of high-level search: hard-to-learn value functions, complex action spaces, presence of dead ends in the environment, or using data collected from diverse experts. We propose a consistent evaluation methodology to achieve meaningful comparisons between methods and reevaluate the state-of-the-art algorithms.
@@ 1 Introduction
Many real-world tasks involve underlying abstract structures that capture their complexity. For example, robotic navigation can be efficiently modeled as a graph search problem, enabling optimal path planning and execution within complex environments [37, 9]. Similarly, sequence alignment in bioinformatics can be effectively modeled as a graph of alignment positions and pathways, facilitating the identification of similarity regions [44, 54]. Hence, addressing these problems in discrete environments, which demand intricate reasoning and include numerous NP-hard problems, is crucial for advancing AI [4].
Tackling these challenges requires efficient planning strategies [5]. A recent, promising approach to planning uses subgoals to decompose complex tasks into sequences of simpler steps, drawing inspiration from how humans plan their actions [30, 21, 33]. Hierarchical planning was successfully applied to tasks with high combinatorial complexity by Subgoal Search (kSubS) [14], followed by other algorithms, including Adaptive Subgoal Search (AdaSubS) [63], Hierarchical Imitation Planning with Search (HIPS) [35], and HIPS-$\varepsilon$ [36]. These methods take high-level steps, which are commonly believed to scale better to longer horizons [7, 38], bypass errors in value approximations [14] and reduce complexity by decomposing the problem into subtasks [56, 63].
While subgoal methods have demonstrated strong results in the literature, our findings reveal that classical, low-level search algorithms can perform at least equally well in certain domains. Thus, it remains an open question in which settings subgoal-based methods should be applied to leverage their advantages.
In this work, we present a comprehensive study exploring how the choice of environments and training data impacts subgoal-based methods for combinatorial reasoning problems. To the best of our knowledge, this is the first systematic study to investigate this relationship. We identify key attributes that enable leveraging the advantage of high-level search: hard-to-learn value functions, complex action spaces, the presence of dead ends in the environment, and the use of data collected from diverse experts. Additionally, our experiments reveal several evaluation pitfalls that are easy to overlook but can lead to misleading conclusions. To address these issues, we propose a consistent evaluation methodology that allows meaningful comparisons between various methods.
@@
  • We identify the key properties of environments and datasets that impact the performance of subgoal-based methods.
  • We analyze the suitability of Hierarchical Reinforcement Learning (HRL) algorithms for combinatorial reasoning and list the necessary attributes.
  • We highlight critical points for reporting reliable results in this domain, such as reporting success rates for the entire search budget.
@@ 2 Related Work
Classical reinforcement learning (RL) algorithms, such as Proximal Policy Optimization (PPO) [49] and Deep Q-Network (DQN) [42], are designed to solve problems modeled as Markov Decision Processes (MDPs) [56]. Recently, there has been a surge in research leveraging large datasets of demonstrations [58, 11, 25, 17] to learn and solve complex real-world decision-making tasks using Imitation Learning (IL) techniques. These methods, including Behavioral Cloning (BC) [55, 48, 58], Inverse Reinforcement Learning [3], and DAgger [47], aim to mimic expert actions provided in datasets.
Subgoal-based methods decompose complex tasks into manageable subgoals, streamlining the solution of complex MDPs. HIRO [43] leverages offline data by relabeling past transitions. HAC [39] learns a hierarchy of high-level policies using hindsight [1]. Graph-based search methods, such as SoRB [18] or DHRL [38], construct a high-level graph of the state space, enabling fast searches for arbitrary goals. Hierarchical Diffuser [7] generates action plans at various timescales using a diffusion model [59]. HIPS [35] employs Policy-guided Heuristic Search [45] to navigate between subgoals generated by VQ-VAE models [57]. HIPS-$\epsilon$ [36] combines high-level and low-level planning to ensure completeness of the search.
Recently, kSubS [14] was introduced as a hierarchical search algorithm for combinatorial domains. To solve the given task, it iteratively constructs a search tree consisting of achievable subgoals. In each iteration, the $k$-subgoal generator predicts states that are $k$ steps closer to the solution, after which the low-level policy computes paths to these subgoals. Nodes are expanded based on a best-first strategy guided by the value function predictions. All components of kSubS are trained on offline datasets using imitation learning.
AdaSubS [63] builds on kSubS by adding multiple subgoal generators trained to output subgoals at various distances. During the planning phase, AdaSubS prioritizes long-distance subgoal generators, retracting to shorter if the search becomes locally stuck. This way, AdaSubS dynamically adjusts its planning strategy based on the difficulty of the current search region.
Classical search algorithms are widely used to solve hard decision-making problems [27, 23, 41, 31, 50, 19, 8]. Best-First Search (BestFS) prioritizes states with the highest predicted values, often using policy to select actions based on top-ranked choices or cumulative confidence thresholds [20]. A* improves upon BestFS by combining value estimates with distances from the root state, enabling more efficient exploration. MCTS [12] builds an asymmetric tree to estimate action values through random simulations, precisely choosing the most promising action. Monte Carlo Tree Search (MCTS) [12] builds an asymmetric tree to estimate action values through random simulations, selecting the most promising actions precisely. Combining MCTS with deep neural networks has led to breakthroughs like AlphaGo and AlphaZero [50, 52].
3 Preliminaries
3.1 Combinatorial environments and algorithmic desiderata
The abstract structure of real-world tasks captures the essence of reasoning. Therefore, our study targets solving discrete combinatorial reasoning environments, a category that includes many NP-hard problems, such as the Traveling Salesman Problem [2], the Knapsack Problem [32], the Rubik’s Cube [53], or the Sudoku [62]. Specifically, in our experiments, we use the Rubik’s Cube, Sokoban, N-Puzzle, and Inequality Theorem Proving (INT) [60]. These are classical benchmarks, explored in various planning works [41, 14], and are known to be NP-hard [15, 13, 46]. Descriptions of these environments are in Appendix A.
The combinatorial reasoning environments are typically characterized by an enormous space of possible configurations, making learning without priors or handcrafted dense rewards prohibitively challenging [^1] For instance, we trained a PPO [49] agent to solve the Rubik’s Cube, but, unsurprisingly, it failed to make any progress due to never reaching the goal in the haystack of $4.3\times 10^{19}$ states, hence never observing a positive reward. This would be the case for any algorithm that learns goal-reaching. Similarly, the space complexity usually limits the number of configurations that can be observed in training. Hence, the solver cannot expect that it will eventually visit a closely located state again, which is essential for methods like [18, 29, 38]. Therefore it follows that the following conditions are necessary for an algorithm to be efficient in solving combinatorial problems:
@@ 3.1 Combinatorial environments and algorithmic desiderata
    Learning from offline demonstrations Finding a solution through exploration is prohibitively hard in combinatorial environments. Thus, the algorithm has to be able to learn from demonstrations.
@@
    Generalization The algorithm will typically visit a given state only once, so it needs to generalize well across the state space.
@@
    Planning The algorithm needs a planning module as it cannot scale in solving NP-hard problems [5] within a fixed computation budget.
@@
Standard hierarchical methods have not been specifically designed for combinatorial problems and, thus, fail to meet the necessary conditions for effective application. For instance, [7, 61] require continuous state or action space, [24] learns from online interactions, [18, 29, 38] assume a good coverage of the whole state space, and [43, 39] do not use planning to determine actions.
3.2 Subgoal methods
Subgoal methods perform planning with high-level steps composed of multiple actions. This high-level guidance may take various forms. We consider subgoal methods characterized by the following components:
  • Subgoal generator that, given a state, outputs subgoals to achieve. That may be a future state, a class of desired outcomes, or any other form of high-level guidance.
@@ 3.2 Subgoal methods
  • Low-level policy that determines low-level paths between subgoals. It may be a trained goal-reaching policy [14, 63], a local search [14, 35], a stored path from previous episodes [18, 38].
  • Planner that determines the order of generating subgoals. Usually, standard planners, e.g., BestFS [14], PHS [35], or their modified forms [63], are repurposed for this goal.
  • Value function that estimates the distance to the solution.
In this work, we will study kSubS and AdaSubS as subgoal methods suitable for combinatorial problems. We also considered HIPS, HIPS-$\varepsilon$, but found them impractical (see Appendix I for details). For reporting, we use the notation kSubS-$n$ to indicate that the subgoal generator was trained for distance $n$. Similarly, AdaSubS-$a$+$b$+$c$ indicates using a set of generators trained for distances $a$, $b$, and $c$.
We compare the performance of subgoal search methods with three widely used low-level algorithms: BestFS, A*, and MCTS. Detailed description, analysis, and pseudocode for each of those low-level algorithms are provided in Appendix F.
@@ 4 Analysis
In this work, we aim to establish a solid empirical foundation for developing subgoal-based search algorithms tailored to combinatorial reasoning problems. To this end, we pose the following fundamental questions:
@@
    Are hierarchical search algorithms indeed proficient in solving combinatorial problems?
@@
    In which environments do hierarchical search algorithms offer an advantage over traditional low-level search strategies? Can we identify specific properties that make an environment amenable to high-level search?
@@
    How should we select the baselines for a fair comparison? What common mistakes should be avoided when reporting experimental results?
To address these questions, we extensively analyze the algorithms and combinatorial reasoning environments. Each experiment is described starting with its motivation, followed by the presentation of our experimental results, and concluding with concise takeaways. [^2] The code used to run the experiments can be accessed via the following link: github.com/subgoal-verification/what-matters-in-hierarchical-search.
Experimental setup
For each experiment, we generate a dataset of approximately $5\times 10^{5}$ solving trajectories. We train the components for each method through imitation learning. Whenever possible, components are shared between algorithms; for instance, all tested algorithms use the same value function. For clarity, we present only the best-performing variants of each method. Tuning the parameters is discussed in Appendix F.
Our primary metric, success rate, is the percentage of solved instances calculated across a set of $500$ instances in a given environment. The success rate depends on the search budget, which is the total number of states visited during the search, including both the subgoals and the low-level steps.
The costs associated with this metric arise from two main sources: generating subgoals and filling low-level paths, each requiring network calls. Therefore, we argue that reporting all search nodes is essential. For instance, a generator might produce a subgoal that is 10 steps ahead of the current state, but reaching this subgoal could be highly costly due to the associated reaching expenses. However, it would be hidden by the other metric only for subgoal methods. A further discussion of this decision is presented in Section 4.3.1.
4.1 Rubik’s Cube
4.1.1 Verifying experimental outcomes
Motivation
Learning to solve the Rubik’s Cube is hard, even for humans. [14, 63] report that both kSubS and AdaSubS can solve it, while the low-level BestFS struggles. We want to verify these claims and understand the reasons behind them.
Experimental Results
Following [63], each method was trained on trajectories obtained by shuffling the cube with $20$ random moves and reversing it. As shown in Figure 2, every tested method solves the environment nearly perfectly. AdaSubS is the most effective when using smaller search budgets, while BestFS takes the lead for larger budgets. A* performs similarly to BestFS, but MCTS is not effective. Thus, using subgoals in this setting is not helpful. This contrasts with the results presented in [14, 63]. Firstly, the low-level algorithms successfully solve the task. Secondly, each subgoal method requires about three times smaller computational budget than reported in [63].
For this study, we reimplemented and trained the components from scratch, with some technical modifications, such as model size, architecture, and hyperparameters, and using up-to-date libraries. We also noticed that BestFS is very sensitive to the quality of its components and hyperparameters (see Appendix F for details). That may explain the substantial difference in performance.

### v3

# UC Berkeley

@@ Abstract
Combinatorial reasoning problems, particularly the notorious NP-hard tasks, remain a significant challenge for AI research. A common approach to addressing them combines search with learned heuristics. Recent methods in this domain utilize hierarchical planning, executing strategies based on subgoals. Our goal is to advance research in this area and establish a solid conceptual and empirical foundation. Specifically, we identify the following key obstacles, whose presence favors the choice of hierarchical search methods: hard-to-learn value functions, complex action spaces, presence of dead ends in the environment, or training data collected from diverse sources. Through in-depth empirical analysis, we establish that hierarchical search methods consistently outperform standard search methods across these dimensions, and we formulate insights for future research. On the practical side, we also propose a consistent evaluation guidelines to enable meaningful comparisons between methods and reassess the state-of-the-art algorithms.
@@ 1 Introduction
Figure 1: Schematic performance comparison of hierarchical methods (AdaSubS, kSubS) and low-level methods ($\rho$-BestFS, $\rho$-A*, $\rho$-MCTS) across six dimensions: handling data collected from diverse sources, learning from clean unimodal demonstrations, avoiding dead ends, performance under high value approximation errors, handling complex action space, and generalizing to out-of-distribution instances.
The ability to solve discrete tasks that require sophisticated reasoning, particularly those involving NP-hard problems, is essential for advancing AI (Bengio et al., 2021). These include complex problems like theorem proving (Wu et al., 2021; Trinh et al., 2024), constraint satisfaction problem (Achiam et al., 2017), molecule alignment (Needleman & Wunsch, 1970; Smith & Waterman, 1981), social network analysis (Kipf & Welling, 2017), or navigation (LaValle, 2006; Choset et al., 2005).
Even driving a car, which typically involves continuous control of steering and speed, requires high-level discrete decision-making, e.g., when to overtake, when to change lanes, or how to navigate through traffic (Kiran et al., 2022).
Addressing that kind of tasks, known as combinatorial reasoning problems, requires efficient planning strategies due to the vast and complex search spaces involved (Bruck & Goodman, 1987). A promising approach to this challenge, inspired by how humans plan their actions (Hull, 1932; Fishbach & Dhar, 2005; Kool & Botvinick, 2014), is hierarchical search. This method breaks down a problem into manageable subproblems, or subgoals, making the overall task more tractable, in contrast to low-level methods that rely on atomic actions for planning. Hierarchical search has been successfully applied to a variety of combinatorial reasoning tasks, as evidenced by methods like Subgoal Search (kSubS) (Czechowski et al., 2021), and further advanced by approaches such as Adaptive Subgoal Search (AdaSubS) (Zawalski et al., 2023), Hierarchical Imitation Planning with Search (HIPS) (Kujanpää et al., 2023a), and HIPS-$\varepsilon$ (Kujanpää et al., 2023b).
Our goal in this paper is to advance research in hierarchical planning and establish a solid conceptual and empirical foundation. We identify four key challenges whose presence highly favors the use of hierarchical search methods: hard-to-learn value functions, complex action spaces, presence of dead ends in the environment, or data collected from diverse sources. Through comprehensive empirical analysis, we demonstrate that hierarchical methods consistently outperform standard search techniques in overcoming these critical obstacles. Furthermore, we propose a consistent evaluation methodology to facilitate meaningful comparisons between methods and reassess current state-of-the-art algorithms. Our findings offer a clearer understanding of when hierarchical approaches should be preferred over low-level methods.
@@
  • We present a comprehensive empirical analysis comparing the performance of hierarchical search methods against low-level search methods across diverse problem settings.
  • We identify problem characteristics that influence performance, providing insights into when hierarchical methods should be favored over low-level methods.
  • We propose a standardized evaluation guidelines that facilitate meaningful and consistent comparisons across different types of search methods.
@@ 2 Related Work
Solving Decision-Making Problems
Decision-making problems are often framed as Markov Decision Processes (MDPs) (Sutton et al., 1999), which can be solved using Reinforcement Learning (RL) algorithms like PPO (Schulman et al., 2017) or DQN (Mnih et al., 2015). These methods learn policies through interaction with the environment. An alternative to learning from trial and error is Imitation Learning (IL), training models directly from offline demonstrations. The availability of large-scale datasets (Walke et al., 2023; Collaboration et al., 2023; Grauman et al., 2022; Dosovitskiy et al., 2017), make it applicable to the most complex domains like robotics (Mandlekar et al., 2018; Edmonds et al., 2017; Kim et al., 2024), autonomous driving (Kelly et al., 2019; Li et al., 2022; Zhang & Cho, 2017), and physics-based control (Kim et al., 2020; Fickinger et al., 2022). Key foundational methods such as Behavioral Cloning (BC) (Sutton & Barto, 1998), Inverse Reinforcement Learning (IRL) (Baker et al., 2009), or DAgger (Ross et al., 2011) have been instrumental in advancing IL for complex environments where direct exploration is less practical. In this work, we use IL to train components for the search methods, such as the policy and value function.
Subgoal Methods
Hierarchical Reinforcement Learning methods tackle complex decision-making tasks by breaking them into subgoals. HIRO (Nachum et al., 2018) reuses past data by goal relabeling. HAC (Levy et al., 2019) builds a multi-layer hierarchy of policies trained with hindsight. Hierarchical Diffuser (Chen et al., 2024) learns to predict future states with diffusion models. Graph-based methods, such as SoRB (Eysenbach et al., 2019) or DHRL (Lee et al., 2022) build a high-level graph of states, which then allow for efficient shortest path finding. GCP (Pertsch et al., 2020) learns to predict middle states between two given observations. Algorithms such as HPG (Ghavamzadeh & Mahadevan, 2003) or H-DDPG (Yang et al., 2018) extend the classical RL algorithms to the hierarchical setting.
In the area of combinatorial reasoning, there has been growing interest in applying HRL techniques. kSubS (Czechowski et al., 2021) introduces a hierarchical search algorithm that iteratively generates subgoals to construct a search tree. Building on this, AdaSubS (Zawalski et al., 2023) incorporates multiple subgoal generators, each trained to predict subgoals at different distances from the target, allowing for dynamic adaptation of the planning horizon based on problem complexity. HIPS (Kujanpää et al., 2023a) and HIPS-$\varepsilon$ (Kujanpää et al., 2023b) perform search using subgoals generated by VQ-VAE models (van den Oord et al., 2017).
Low-level Search Algorithms
Traditional search algorithms like Best-First Search (BestFS), A* (Cormen et al., 2009; Russell & Norvig, 2009), and Monte Carlo Tree Search (MCTS) (Veness et al., 2009; James et al., 2017) have long been the foundation for solving complex decision-making problems. Recent advancements have improved these methods by integrating neural network-based heuristics, improving their efficiency in large search spaces (Silver et al., 2018; Yonetani et al., 2021). A variant of $\rho$-BestFS used in (Czechowski et al., 2021; Zawalski et al., 2023), leverage heuristics learned through behavioral cloning to guide search. More recent algorithms, like PHS (Orseau & Lelis, 2021) or LevinTS (Orseau et al., 2023), combine policy-driven and value-based approaches, offering both theoretical guarantees and strong empirical performance. Additionally, PDDL planners (Haslum et al., 2019) solve decision-making problems by using predefined action models and goals, with domain-independent planners offering broad applicability, while domain-specific ones achieve higher performance in specialized tasks.
Empirical Studies on Algorithmic Performance
Our work aligns with recent empirical studies that investigate the conditions under which various algorithmic approaches excel. For instance, (Andrychowicz et al., 2020) investigates how specific design choices influence the performance of PPO, while other research compares offline reinforcement learning with behavioral cloning (Kumar et al., 2022) or explores design choices for language-conditioned robotic imitation learning (Mees et al., 2022). In this paper, we focus on hierarchical search in combinatorial reasoning problems, specifically studying the conditions where hierarchical methods outperform low-level planners. To the best of our knowledge, this is the first systematic study of the relationship between hierarchical and low-level search in this context.
3 Combinatorial Environments
Our study targets solving combinatorial environments – domains in which the number of possible configurations or decisions grows exponentially with the problem size, making them highly challenging to solve. This class includes several NP-hard problems, such as the Traveling Salesman Problem (Applegate et al., 2006), the Rubik’s Cube (Singmaster, 1981), Sokoban (Culberson, 1997), or solving non-linear inequalities (Sahni, 1974). To efficiently solve combinatorial problems an algorithm should have the following key properties:
@@ 3 Combinatorial Environments
    Learning from offline data. Since combinatorial reasoning environments are characterized by a large space of possible configurations, learning without priors or handcrafted dense rewards is infeasible [^1] For instance, we tested PPO (Schulman et al., 2017) on the Rubik’s Cube, but, unsurprisingly, it failed to make any progress due to never reaching the goal in the haystack of $4.3\times 10^{19}$ states, hence never observing a positive reward. Thus, the algorithm has to be able to learn from additional offline data, such as demonstrations.
@@
    Combinatorial space abstraction. The space complexity significantly restricts the fraction of observable states. As a result, it is unrealistic to expect repeated visits to nearby states, an assumption that some approaches implicitly rely on.
@@
    Planning. Methods that don’t use search and follow a single action trajectory are inherently limited by computational complexity, since they can perform only a constant number of operations before choosing an action. Solving NP-hard problems within a fixed computation budget is computationally infeasible (Bruck & Goodman, 1987).
@@
Many hierarchical methods have not been designed for combinatorial problems, so they fail to meet the listed conditions and cannot be expected to be efficient in these applications. For instance, (Chen et al., 2024; Yang et al., 2018) require continuous state or action space, (Ghavamzadeh & Mahadevan, 2003) learns only from online interactions, (Eysenbach et al., 2019; Huang et al., 2019; Lee et al., 2022) assume a good coverage of the whole state space, and (Nachum et al., 2018; Levy et al., 2019) do not use planning to determine actions.
4 Subgoal Methods
Subgoal methods, or hierarchical methods, are a family of algorithms designed to solve complex decision-making tasks by breaking down the overall objective into smaller, more manageable subgoals (Sutton et al., 1999). Instead of searching for a sequence of low-level actions that directly lead from the initial state to the goal, the agent first identifies high-level intermediate targets – subgoals – that guide the trajectory toward the final goal. The use of subgoals is widely considered as a method that scales better to longer horizons (Chen et al., 2024; Lee et al., 2022), mitigates errors in value approximations (Czechowski et al., 2021), and reduces overall complexity by decomposing the problem into smaller subproblems (Sutton et al., 1999; Zawalski et al., 2023). The process of searching involves the following components:
  • Subgoal generator that, given a state within the search tree, outputs subgoals to be achieved. For instance, a subgoal may be a future state (Czechowski et al., 2021; Zawalski et al., 2023) or a class of desired outcomes (Jiang et al., 2019; Panov & Skrynnik, 2018). The generator is used by the planner to construct a search tree of subgoals.
@@ 4 Subgoal Methods
  • Low-level policy that determines a path of low-level actions between subgoals. For instance, it may be a trained goal-reaching policy (Czechowski et al., 2021; Zawalski et al., 2023), a local search (Czechowski et al., 2021; Kujanpää et al., 2023a), or a stored path from previous episodes (Eysenbach et al., 2019; Lee et al., 2022).
  • Planner that determines the order in which subgoals are generated. Standard planning algorithms like BestFS (Czechowski et al., 2021), PHS (Kujanpää et al., 2023a), or their modified forms (Zawalski et al., 2023), are typically used.
  • Value function that estimates the distance between the given state and the goal state. The planner uses this information to select the next node to expand with the subgoal generator. In some works it is also called heuristic value.
In our experiments, we use kSubS Czechowski et al. (2021) and AdaSubS Zawalski et al. (2023) as subgoal methods well-suited for combinatorial problems, as they satisfy the conditions formulated in Section 3. We also experimented with HIPS and HIPS-$\varepsilon$ (Kujanpää et al., 2023a; b), but these methods generally fail to solve the problems within a reasonable computational budget. Therefore, their results are omitted from the main text and discussed in see Appendix I.
We compare the performance of the selected subgoal approaches against three popular low-level methods: BestFS, A*, and MCTS. To ensure a fair comparison and improve efficiency, we augment these algorithms by using a trained policy to select the top actions before each node expansion. We refer to them as $\rho$-BestFS, $\rho$-A*, and $\rho$-MCTS. A detailed description, analysis, and pseudocode for each of these algorithms can be found in Appendix F. See also Appendix H for diagrams explaining different search methods.
4.1 Training Components
In our experiments, the models for both subgoal methods and low-level searches were trained using imitation learning, following standard practice (Nair et al., 2018; Czechowski et al., 2021). Specifically, we collected a dataset of approximately 500 000500000500\,000500 000 trajectories for each environment. Trajectories are sequences of consecutive states and actions leading to the goal state. We used various methods of dataset collection, like hand-crafted algorithms, trained policies, reversed random shuffles, and others, which let us to study the influence of training data characteristics on the performance of search methods.
To ensure a fair comparison, all methods shared common components whenever applicable (e.g., each method uses the same value function). This allows us to focus on the differences between the search algorithms, rather than heuristic biases. No additional heuristics were used, ensuring that performance differences arise solely from the algorithmic approaches.
More details on training the components, including specific objectives, are provided in Appendix D.
4.2 Performance Metric
Our performance metric is the success rate, defined as the percentage of problem instances solved within a given complete search budget. The complete search budget is the total number of visited states in the search tree. In particular, for subgoal methods, the budget includes both the generated subgoals and the states visited by the low-level policy used to connect these subgoals.
By accounting for the total number of visited states, this metric provides a unified and fair comparison of search efficiency across different methods. We argue that reporting only the number of visited subgoal nodes would unfairly favor subgoal methods (see Appendix I for details).
@@ 5 Analysis
We investigate how environmental properties and training data influence the performance of hierarchical methods compared to low-level search approaches in combinatorial reasoning tasks. While previous works (Czechowski et al., 2021; Zawalski et al., 2023; Kujanpää et al., 2023a; b) show a considerable advantage of hierarchical methods, our experiments reveal that this advantage is not consistent across all scenarios (see Figures 6 or 6 for specific examples). Specifically, we answer the following research questions:
@@
    Is hierarchical search more effective than low-level search for solving combinatorial reasoning problems?
@@
    What environmental properties and characteristics of the training data amplify performance differences? When hierarchical search should be preferred over low-level search?
@@
    What pitfalls should be avoided when interpreting experimental results?
To address these questions, we conducted a wide range of experiments comparing subgoal and low-level search algorithms across a variety of combinatorial reasoning tasks. Below, in each subsection we summarize the key findings that reveal the most significant factors affecting performance, followed by a brief discussion. For each finding, we link it to the relevant research questions. The extended analysis of these factors can be found in Appendix B.
We present our findings using the Rubik’s Cube, Sokoban, N-Puzzle, and Inequality Theorem Proving (INT) (Wu et al., 2021) environments. These classical benchmarks are widely used in planning research (McAleer et al., 2019; Czechowski et al., 2021) and are known to be NP-hard (Demaine et al., 2018; Culberson, 1997; Ratner & Warmuth, 1986). Detailed descriptions of these environments can be found in Appendix A.
All methods in our study were trained using imitation learning, with each approach sharing the same value function, as outlined in Section 4.1. To ensure fair comparisons, we measured complete search budgets, in contrast to counting only high-level search nodes, to avoid giving any unfair advantage to subgoal methods, as discussed in Section 4.2 (which contributes to the research question Q3).
5.1 Subgoal Methods Benefit from Diverse Sources of Data
Achieving superhuman performance in complex tasks often involves large-scale datasets of demonstrations obtained from agents with varying skill levels and strategies (Silver et al., 2016). By training models on data collected from a variety of solvers and testing them in the Rubik’s Cube and N-Puzzle environments, we show that the variability in training data has a significant impact on the performance of search algorithms.
Figure 2: Solving the Rubik’s Cube. Components are trained on data from 4 different solvers.
Figure 3: Solving the N-Puzzle. Components are trained on data from 2 different solvers.
Figure 6: Solving N-Puzzle. Components are trained on an algorithmic solver.
As shown in Figures 3-3, subgoal methods consistently outperform low-level methods by a wide margin (Q1). However, when the training dataset is limited to a single source of demonstrations – whether the demonstrations are long and structured or short and direct – this performance gap disappears (see Figures 6-6). Notably, subgoal methods, particularly AdaSubS, maintain stable performance across all training setups, while low-level methods are highly sensitive to the characteristics of the training data.
To explain those results, we found that value functions trained on diverse data often fail to assign consistently low values to the initial states of tasks. When demonstrations differ significantly in their length or execution style, the value function learns this variation, leading to inconsistent value predictions. Hierarchical methods can overcome this issue by relying on subgoals. Subgoals enable the agent to make long steps toward the solution, effectively bypassing regions of the state space where the value function is inconsistent or noisy, as it does not need to assess every small step along the way (this property is further studied in Section 5.2). In contrast, low-level methods operate on a finer, step-by-step level, executing small, atomic actions. This makes them more sensitive to the variability in the value function because they must evaluate each intermediate state on the way.
More detailed analysis of the experiments involving diverse data sources is provided in Appendix B.1.